FAX2PS(1) FAX2PS(1) NAME fax2ps−convertaTIFFfacsimiletocompressedPOSTSCRIPTTM SYNOPSIS fax2ps[options][file...] DESCRIPTION fax2psreadsoneormoreTIFFfacsimileimagefilesandprintsacompressedformofPOSTSCRIPT onthestandardoutputthatissuitableforprinting. Bydefault,eachpageisscaledtoreflecttheimagedimensionsandresolutionsstoredinthefile. The−xand−yoptionscanbeusedtospecifythehorizontalandverticalimageresolutions (lines/inch),respectively.Ifthe−Soptionisspecified,eachpageisscaledtofillanoutputpage. Thedefaultoutputpageis8.5by11inches.Alternatepagedimensionscanbespecifiedininches withthe−Wand−Hoptions. Bydefaultfax2psgeneratesPOSTSCRIPTforallpagesinthefile.The−poptioncanbeusedto selectoneormorepagesfromamulti-pagedocument. fax2psgeneratesacompressedformofPOSTSCRIPTthatisoptimizedforsendingpagesoftextto aPOSTSCRIPTprinterattachedtoahostthroughalow-speedlink(suchasaserialline).Each outputpageisfilledwithwhiteandthenonlytheblackareasaredrawn.ThePOSTSCRIPTspecificationoftheblackdrawingoperationsisoptimizedbyusingaspecialfontthatencodesthe move-drawoperationsrequiredtofilltheblackregionsonthepage.Thiscompressionscheme typicallyresultsinasubstantiallyreducedPOSTSCRIPTdescription,relativetothestraightforward imagingofthepagewithaPOSTSCRIPTimageoperator.Thisalgorithmcan,however,beineffectiveforcontinuous-toneandwhite-on-blackimages.Fortheseimages,itsometimesismoreeffi-cienttosendtherasterbitmapimagedirectly;seetiff2ps(1). OPTIONS −pnumber Printonlytheindicatedpage.Multiplepagesmaybeprintedbyspecifyingthisoption morethanonce. −xresolution Useresolutionasthehorizontalresolution,indots/inch,oftheimagedata.Bydefault thisvalueistakenfromthefile. −yresolution Useresolutionastheverticalresolution,inlines/inch,oftheimagedata.Bydefault thisvalueistakenfromthefile. −SScaleeachpageofimagedatatofilltheoutputpagedimensions.Bydefaultimages arepresentedaccordingtothedimensioninformationrecordedintheTIFFfile. −Wwidth Usewidthasthewidth,ininches,oftheoutputpage.Thedefaultpagewidthis8.5 inches. −Hheight Useheightastheheight,ininches,oftheoutputpage.Thedefaultpageheightis11 inches. DIAGNOSTICS SomemessagesaboutmalformedTIFFimagescomefromtheTIFFlibrary. Variousmessagesaboutbadlyformattedfacsimileimagesmaybegeneratedduetotransmission errorsinreceivedfacsimile.fax2psattemptstorecoverfromsuchdataerrorsbyresynchronizing decodingattheendofthecurrentscanline.Thiscanresultinlonghorizontalblacklinesinthe resultantPOSTSCRIPTimage. March16,1995 1 FAX2PS(1) FAX2PS(1) NOTES IfthedestinationprintersupportsPOSTSCRIPTLevelIIthenitisalwaysfastertojustsendthe encodedbitmapgeneratedbythetiff2ps(1)program. BUGS fax2psshouldprobablyfigureoutwhenitisdoingapoorjobofcompressingtheoutputandjust generatePOSTSCRIPTtoimagethebitmaprasterinstead. SEEALSO tiff2ps(1),libtiff(3) March16,1995 2